OpenStudy (anonymous):

the volume of a sphere is found using the formula v=4/3πr^3 where r is the radius. what is the volume if the radius is 3x^3 a. 36πx^9 b. 36πx^6 c. 12πx^9 d. 12πx^6

OpenStudy (kamibug):

We just gotta plug in the value of r into the formula and simplify. :) \[V = \frac{ 4 }{ 3 }\left( 3x ^{3} \right)^{3}\]

OpenStudy (kamibug):

\[V = \frac{ 4 }{ 3 }\left( 9x ^{9} \right)\]

OpenStudy (kamibug):

Oops, I just noticed I've been missing the pi sign, Lol. :P Let's include it this time! xD \[V = 12\pi x^9\]

OpenStudy (anonymous):

yeah im lost ahaha

OpenStudy (kamibug):

Well first we put 3x^3 where r is in the formula because we are told that r = 3x^3. Then we had to cube the value 3x^3 because the formula asks for r^3. And so when you raise an exponent to the power of another exponent, you need to multiply them. The two exponents are 3 so 3*3=9. And the coefficient of x also had to be cubed since it was in the parenthesis. 3^3=9. So we got 9x^9. And so the formula turned into 4/3pi9x^9. The last step was to multiply 4/3 and 9. 4/3 * 9 = 12. So that left us with 12pix^9.

OpenStudy (kamibug):

Does this help somehow? I know it's super long Lol sorry. :)

OpenStudy (anonymous):

that makes sense

OpenStudy (kamibug):

Okiedokey, so you've got your answer now! ~~~(^_^)~~~

OpenStudy (anonymous):

thanks (: